solve the simultaneous equation x+2y=12 and 3x-5y=3
1
Expert's answer
2013-03-05T09:52:50-0500
x+2y=12 and 3x-5y=3
x=12-2y from the first then we substitute obtained expression into the second equation 3*(12-2y)-5y=336-6y-5y=3
-11y=-33
y=3 then x=12-2*3=6
So the answer is x=6,y=3